A mixture of gases at a pressure 7.50 x 104 Nm-2 has the volume composition 40% N2; 35% O2; 25% CO2. (a) What is the partial pressure of each gas? [9] (b) Deduce the partial pressure of the nitrogen and oxygen if the carbon dioxide is removed by the introduction of some sodium hydroxide pellets? 


1
Expert's answer
2021-04-28T02:46:20-0400

Partial pressure = mole fraction × total Pressure

N2 = 0.4× 7.5 × 104 Nm-2

= 3 × 104 Nm-2

O2 = 0.35 × 7.5 × 104 Nm-2

= 2.625 × 104 Nm-2

CO2 = 0.25 × 7.5 × 104 Nm-2

= 1.875 × 104 Nm-2

N2 and O2

= 0.75 × 7.5 × 104 Nm-2

= 5.625 × 104 Nm-2